I need exact solutions to three decimal places for the following:
(x-6)^2=1

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

I need exact solutions to three decimal places for the following:
(x-6)^2=1
----------
(x-6)^2-1 = 0
Factor:
(x-6-1)(x-6+1) = 0
(x-7)(x-5) = 0
x = 7.000 or x = 5.000
Cheers,
Stan H.

Answer by Alan3354(69443)

(x-6)^2=1
x-6 = ±1
x = 5, -7